A phcn service is made up of seven strands of aluminum wire each wire has a diameter of 1.7mm calculate the cross sectional area of the service cable

1
Expert's answer
2022-03-28T14:24:34-0400

n=7

"r=0.85mm"

Area

"A=n\pi r^2"


"A=7\times3.14\times(0.85\times10^{-3})^2=1.588\times10^{-5}m^2"